# Signal attenuation enables scalable decentralized multi-agent reinforcement learning over networks

### v1

@@ V Conclusion
In this work, we have considered the specific use-case of power allocation for target detection in a radar network to illustrate how signal strength decay properties inherent in wireless communications and radar networks can be used to develop decentralized, scalable MARL methods. Limitations of this work include the action-independent nature of our dynamics model, the lack of value function approximation in our algorithms, and the lack of experimental validation of our approach. Future directions include performing convergence analysis of Algorithms 1 and 2, developing practical variants of our algorithms leveraging value function approximation, extension of our approach to additional applications in communications and radar networks, and extension of our approach to joint motion planning and power allocation in target detection in mobile radar networks.

### v2

@@ V Conclusion
In this work, we have considered the specific use-case of power allocation for target detection in a radar network to illustrate how signal strength decay properties inherent in wireless communications and radar networks can be used to develop decentralized, scalable MARL methods. Future directions include convergence analysis of Algorithms 1 and 2, experimental evaluation of neural network-based versions of our algorithms, extension of our approach to additional applications in communications and radar networks, and extension of our approach to joint motion planning and power allocation in target detection in mobile radar networks.